[Brahms-l] Shift report 20040305 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Fri Mar 05 2004 - 08:01:30 EST
------------------------------------------------------------
Date: Fri Mar  5 08:01:29 2004, Shift: 00:00-08:00
Supervisor: kh, with: 
Log book#15 pages 75 - 75

Shift Summary:
------------------------------------------------------------
1. Beam from 4:20 on
   Total 3:40 hrs
   
2.  FS 8deg.@B3450   MRS 90deg@B1050 
   run     tr2    tr3
  10765    96K    103K
  10766    27k     40k
    FS 8 deg @B1723  MRS 90deg@B1050
  10767    21k     15k    trig2 = ZDC + H1 + H2
                               (trfs removed from trig2)
  10778    41k     30k    trig2 = ZDC + H1 + H2 
                               (trfs removed from trig2)
  10779      in progress  trig2 restored to nominal, 
                           here only trig4+5+6


3. 
  a)
   Investigated why magtrip segfaults and discovered 
   that data from the Keithley units has not been written 
   since 16-Jan.  I don't know why.  The
   monitorKeithley.perl routine seems to be running on 
   opus.  There does not seem to be a lack of disk space 
   on pii3 where the DB is stored.  And the rest of the 
   RUNDB seems to be working.  We need an expert to 
   investigate.

  b)
   The DAQ suddenly stopped taking data.  I traced it to 
   tpc2 (and perhaps mtpc2 as well)  After trying various 
   suggestions from the problem FAQ, cycling the power on 
   the TPC remote interface made it start working again.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Mar 5 08:01:43 2004

This archive was generated by hypermail 2.1.8 : Fri Mar 05 2004 - 08:02:04 EST